OpenStudy (anonymous):

what is 3 7/10 - (-1 1/5)=

OpenStudy (anonymous):

@ItsKoreyyy

OpenStudy (denonakavro):

Hey can you be more clear writing the equation

OpenStudy (anonymous):

Give me a minute, this could take a few minutes to type out and solve.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ 37 }{ 10}-(-1\frac{ 1 }{ 5 })\] \[\frac{ 37 }{ 10} - (-\frac{ 6 }{ 5 })\] \[\frac{ 37 }{ 10} - \frac{ -6 }{ 5 }\] \[\frac{ 49 }{ 10 }\] = 4\[\frac{ 9 }{10 }\] is your complete answer.

OpenStudy (anonymous):

Wait it inserted wrong with the answer one sec

OpenStudy (anonymous):

\[4\frac{ 9 }{ 10 }\] This is your answer :)

OpenStudy (anonymous):

thx